# Multimodal Foundation Models for Early Disease Detection

### v1

@@ VII Conclusion
We outline a framework for multimodal foundation models aimed at healthcare diagnostics. The idea is simple: instead of relying on one type of data, such as EHRs, the model brings together records, images, genetic information, and signals from wearable devices. A transformer-based setup is used to handle this mix. Each type of input is first processed by its own encoder, so important details are not lost. These are then linked through attention mechanisms, and pretraining helps the model build general structure before it is tuned for specific clinical tasks. The design also stresses aspects that are often overlooked, like data governance, model monitoring, and interpretability. In practice, these matter as much as raw accuracy, since clinicians need to understand decisions, trust outputs, and know that data quality is maintained. The framework structure is meant to be flexible across diseases and patient groups. Over time, we hope this kind of system could allow earlier diagnosis, better risk prediction, and treatment tailored to individuals—steps toward precision medicine.

### v2

@@ VII Conclusion
We outline a framework for multimodal foundation models designed for healthcare diagnostics. In practice, clinical datasets remain fragmented, frequently include missing modalities, and carry high-dimensional inputs that complicate integration and training [19]. Our core idea is to move beyond reliance on a single or fragmented data source, and instead integrate records, images, genomic profiles, and wearable signals into a unified representation. A transformer-based setup is used to handle this mix. Each type of input is first processed by its own encoder, so important details are not lost. These are then linked through attention mechanisms, and pretraining helps the model build general structure before it is tuned for specific clinical tasks. The design also stresses aspects that are often overlooked, like data governance, model monitoring, and interpretability. The framework structure is meant to be flexible across diseases and patient groups. Over time, we hope this kind of system could allow earlier diagnosis, better risk prediction, and treatment tailored to individuals—steps toward precision medicine.